OpenStudy (anonymous):

Enter the equation of the circle described below. Center (3, -2), radius = 5

OpenStudy (solomonzelman):

\(\large\color{slate}{ (x-\color{red}{h})^2+(y-\color{blue}{k})^2=\color{green}{r}^2 }\) is the equation of a circle, where: \(\large\color{slate}{ (\color{red}{h},\color{blue}{k}) }\) is the center \(\large\color{slate}{ \color{green}{r} }\) is the radius

OpenStudy (anonymous):

Equation of a circle = \[(x-h)+(y-k)^2 = r^2\] where your center is (h, k) and r is the radius.

OpenStudy (anonymous):

Well Solomon is much faster and more colourful :P

OpenStudy (solomonzelman):

I started before you,

OpenStudy (solomonzelman):

but anyways, helpzzz, is this circle-equation makins sense?

OpenStudy (anonymous):

Ah, my bad then!

OpenStudy (solomonzelman):

why is it your bad?

OpenStudy (solomonzelman):

everyone has a right to help, if you are ging on that.

OpenStudy (anonymous):

Haha, thanks :). Does this make sense? :P

OpenStudy (solomonzelman):

@hhelpplzzzz say something, please:)